OpenStudy (anonymous):

solve by factoring. 8x^2+6x+1=0

OpenStudy (anonymous):

(4x + 1) ( 2x +1 ) = 0 x = -1/4 and x = -1/2

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

no problem :)